OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (dangerousjesse):

It looks null..

OpenStudy (anonymous):

for every x in f(x) replace with the function g(x)

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

-40x - -24 ?

OpenStudy (anonymous):

the -40x is correct, but it's not 24

OpenStudy (anonymous):

you forgot to add the 7 i guess

OpenStudy (anonymous):

f(g(x)) = -4(10x - 6) Distribute the -4 and you get -40x and -4 x 6 = 24...I don't see understand?

OpenStudy (anonymous):

f(x) = -4x + 7 g(x) = 10x - 6 Find f(g(x)) When I evaluate expressions that include multiple parentheses, I look for the first right parenthesis, then I back up to the prior left parenthesis and look at what I see. The value for g(x) is what we will put into the first term Change g(x) in your FIND statement to 10x-6 Find f(g(x)) Find f(10x-6) Don't let the fact that "x" is now equal to 10x-6 confuse you. Where you see f(x) = -4x + 7 we put the entire (10x-6) where the x is f(x) = -4(10x-6) +7 -4 * 10x = -40x -4 * -6 = +24 f(x) = (-___ + ___) + 7 You finish it from here.

OpenStudy (anonymous):

f(g(x)) = -4(10x - 6)\(\color{red}{+7}\)

OpenStudy (anonymous):

your method is correct but you forgot the 7 as i thought

OpenStudy (anonymous):

I'm confused as to what you'd add the 7 to...

OpenStudy (anonymous):

f(g(x))= -4(10x - 6)+7. just substitute g(x) value in f(x) expression.

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

you can add a number only to other numbers, so the 7 is added to the 24 to make 31